# GAPS: Guiding Dynamic Android Analysis with Static Path Synthesis

### v1

@@ Abstract.
Dynamically resolving method reachability in Android applications remains a critical and largely unsolved problem. Despite notable advancements in GUI testing and static call graph construction, current tools are insufficient for reliably driving execution toward specific target methods, especially those not embedded in a graphical component (e.g., libraries’ methods), a capability essential for tasks such as vulnerability validation, debugging, and behavioral analysis.
We present GAPS (Graph-based Automated Path Synthesizer), the first system that integrates static, method-guided call graph analysis with dynamic, interaction-driven execution. GAPS performs a lightweight backward traversal of the call graph, guided by data-flow analysis, to reconstruct paths reaching the target methods. These paths are then translated into instructions that guide runtime app exploration.
On the AndroTest benchmark, GAPS statically identifies paths to reach 88.24% of the target methods in just 4.27 seconds per app and dynamically reaches 57.44% of them. In contrast, state-of-the-art dynamic interaction tools show significantly lower reachability over three runs: APE, one of the best model-based GUI testers, achieves 12.82%, while GoalExplorer, a hybrid analysis tool, reaches 9.69%, and Guardian, an LLM-based UI automator, reaches 17.12%. Static analysis tools also fall short: FlowDroid and DroidReach identify paths to reach 58.81% and 9.48% of the targets, requiring 35.06 seconds and 23.46 seconds per app, respectively.
Finally, an evaluation on the 50 most downloaded real-world apps demonstrates GAPS’s practical utility in analyzing security-critical code under a realistic scenario. With an average static analysis time of 278.9 seconds, GAPS statically reconstructs paths to 62.03% of the target methods and dynamically reaches 59.86% of them.
@@ 1. Introduction
Android dynamic analysis is a process aimed at inspecting an Android application’s behavior during its runtime execution. It can be used to verify security vulnerabilities, measure performance, and detect malicious activities. Complementary to static analysis, which inspects the app without executing it, dynamic analysis involves running the app in a controlled environment, such as a real device, an emulator, or a sandbox, while monitoring its interactions with the system, network, and other components. Even though some apps, such as the Android Operating System (OS) services, can come without a Graphical User Interface (GUI), most apps require interaction with the available GUI components to change their status. Thus, dynamic analysis for Android apps mainly includes GUI testing tools. This choice already comes with an important limitation: such tools are able to reach only code within a GUI component and not any code inside an app. In fact, proving the reachability of an app target method is still an open research question (browserprivacy; monkeylimitations; androtest).
@@
The evolution of Android GUI testing tools has gone through different phases, each one leveraging an incrementally sophisticated approach. The very first ones were pseudo-random input generators, such as Monkey (monkey), which lack of understanding of the app semantics and of the UI hierarchy, thus being inefficient for deep functional testing, but very useful for finding shallow random crashes in an app. Then, GUI testers started embedding models, such as finite state machines, to better reproduce common interaction scenarios and maximize exploration. Notably, APE (ape) is a model-guided GUI testing tool that has been recognized in prior studies for its effectiveness in improving code coverage and detecting GUI-related bugs (wang2021infrastructure; su2021benchmarking; akinotcho2024mobile). A hybrid solution, GoalExplorer (goalexplorer), builds a Screen Transition Graph (STG) to statically map relationships between activities and then guides automatic interaction using Stoat (stoat). The advent of machine learning and deep-learning techniques paved the way for GUI testers to autonomously discover new test paths following the guidance of optimization algorithms (ares). Finally, the introduction of Large Language Models (LLMs) allowed GUI testers to leverage semantic and contextual information to solve tasks closer to those performed by humans. Among recent advancements, Guardian (guardian) stands out for leveraging LLMs to guide UI exploration, showing competitive results across task completion benchmarks. Despite the multiple contributions in the field of Android automatic interaction (columbus; androtest; stochtesting; sapienz; timemachine; dynodroid), the efforts made so far have been mainly driven by either increasing the app code coverage or by replacing a real user in the completion of a task (e.g., “add an event to the calendar app”).
In this paper, we present our intuition that allows to finally guarantee method level reachability: the directed exploration of an app can be performed by following a path that connects an app entry point to the target code. Such path can be statically computed by traversing the app call graph, which shows the calling relationships between methods in the app.
@@
At first, we evaluated the performance (i.e., target method reachability) of the best state-of-the-art tools designed for Android GUI interaction and call graph reconstruction. In particular, we randomly selected 50 target methods in each app among the 56 open-source ones [^1] https://github.com/samudoria/GAPS/tree/main/evaluation_data/androtest_results/ of the AndroTest (androtest) dataset, a standard benchmark in the field of Android automatic interaction, still used by current state-of-art (columbus; androtest; stochtesting; sapienz; timemachine; dynodroid; akinotcho2024mobile) . We, then, tested the performance of the selected tools. We started with the GUI testers, evaluating their ability to dynamically reach the target methods within a 5-minute timeout for each app, over three runs: on average, Guardian (guardian) was able to reach only 17.12% of the target methods, while APE (ape) and GoalExplorer (goalexplorer) achieved an average reachability rate of just 12.82% and 9.69% respectively, within the same time limit. In terms of call graph reconstruction, we chose FlowDroid (flowdroid), the best static analysis tool for such purpose (cg_soundness), and DroidReach (droidreach), the only available static path reconstruction tool: FlowDroid was able to reach 58.81% of the target methods, DroidReach was only able to reach 9.48% of the methods. Our findings confirm the lack of both static and dynamic solutions for reaching a target method in an app.
@@
To this end, we design GAPS (Graph-based Automated Path Synthesizer), a framework that combines static analysis with guided dynamic execution to reach target methods in Android apps. Statically, GAPS performs target-oriented, context-sensitive call graph construction via backward traversal, rather than whole-program analysis, restricting its analysis to the subset of the program relevant to a given target method. It uses data-flow analysis techniques, such as points-to-analysis and constant propagation, to statically reconstruct partial execution paths using backward traversal from a target method, resolve conditional statements, and identify triggering UI elements. These results are compiled into high-level interaction plans that dynamically drive the app to execute the desired behavior.
@@
Additionally, we evaluated GAPS on a real-world dataset consisting of the 50 most downloaded applications on the Google Play Store. This experiment serves as a use case scenario to demonstrate how GAPS can be employed in a realistic scenario to assess the reachability of security-relevant code. We began by running SPECK (speck), a rule-based static analyzer, to identify potential vulnerabilities and their corresponding method locations. These methods were then used as targets for GAPS to assess their reachability through both static and dynamic analyses. Our findings show that the static analysis module completed in an average of 278.9 seconds and was able to statically reach 62.03% of the target methods. The dynamic analysis achieved 59.86% of reachability.
@@
  • We design GAPS, a novel approach that integrates static and dynamic analyses to enable method-level reachability in Android apps. We release the framework and the evaluation data and setup in open source [^2] https://github.com/samudoria/GAPS/.
  • We evaluate GAPS against AndroTest, the state-of-the-art dataset for automatic interaction of Android apps. GAPS is able to generate at least one path to reach 88.24% of the target methods in 4.27 seconds on average, and to automatically reach 57.44% of them while never exceeding the 5-minute timeout.
  • We use GAPS in a real use case scenario, which is proving the reachability of potentially vulnerable methods in the 50 most downloaded real-world apps from the Google Play Store. With an average static analysis time of 278.9 seconds  we were able to statically reach 62.03% of the methods and achieve a runtime reachability of 59.86%.
@@ 2.1. GUI-Based Test Input Generators
Dynamic test input generators, often called GUI testers, explore Android apps by interacting with their GUIs. Their core functionality relies on repeatedly querying the device for the current view hierarchy, which is retrieved in XML format and lists all the interactable UI elements on the screen, such as buttons, checkboxes, text fields, and menus. Each GUI tester uses this structured representation to decide which element to interact with next. This decision process is guided by a strategy that may be pseudo-random (e.g., Monkey (monkey)), model-based (e.g., APE (ape)), machine-learning based (e.g., ARES (ares)), or semantic (e.g., Guardian (guardian)).
GUI testers are typically designed to maximize code coverage or detect UI-related bugs. Pseudo-random strategies rapidly issue events for lightweight testing, while model-based and machine-learning based tools interact heuristically with the UI, often guided by internal models of app states or by optimization algorithms, and prioritize breadth and novelty of interactions rather than reaching specific behaviors.
@@
Despite these advances, GUI testers, whether coverage-oriented or goal-driven, are fundamentally limited in their ability to consistently and specifically reach methods in an app, which results in an inefficient exploration. Moreover, they are definitely not able to reach methods that are not embedded in a GUI component (e.g., a library method).
@@ 2.2. Call Graph Generation
Several tools, such as FlowDroid (flowdroid) and Amandroid (amandroid), attempt to build inter-procedural call graphs using context- and flow-sensitive analysis. These tools are computationally intensive and face scalability challenges due to their exhaustive exploration of all app edges. For instance, a previous study (mudflow) reported that even with 730GB of RAM and 24 hours of analysis time, FlowDroid failed to analyze certain applications. Additionally, Android call graphs are known to suffer from unsoundness, with missing or incorrect edges caused by unresolved dynamic behavior (cg_soundness). More specifically, many state-changing events for an Android app are not declared in its code, but are instead part of the Android framework, which implicitly invokes callback methods, such as onCreate, onClick (edgeminer; columbus). In addition, Android apps consist of multiple interacting components (i.e., Activities, Services, BroadcastReceivers), each governed by its own lifecycle. Accurately modeling these lifecycles and their transitions is critical for generating meaningful call graphs, but also difficult since they involve implicit invocations and asynchronous handlers to statically resolve. ICC introduces further complexity, as components interact by sending Intent messages that dynamically determine control-flow across components. Without proper modeling of ICC, the resulting call graph is incomplete and may miss valid transitions altogether (iccta). Finally, static analysis must also consider conditional logic. Even when a path exists between an entry point and a method, it may be guarded by a conditional statement.
Altogether, these challenges make it clear that full call graph construction for Android apps is not only resource-intensive, but also fundamentally limited in soundness and precision (cg_soundness). Moreover, complete call graph construction is not always necessary or even desirable when the analysis goal is to verify the reachability of a specific method. A full call graph introduces the risk of path explosion, especially in large apps with deeply nested lifecycles, asynchronous flows, and extensive use of libraries.
@@ 3. Related Work
In this section, we discuss current state-of-the-art research related to GUI testers, call graph generation and hybrid analysis.
@@ 3.1. GUI Testers
Automated GUI testing for Android has been approached with various techniques, but generally shares the goals of maximizing code coverage and finding UI bugs. Android’s built-in Monkey tool (monkey) is the archetype of random GUI testing. It simply issues pseudo-random UI and system events with minimal configuration, making it extremely easy to use. Variants like SmartMonkey (smartmonkey) and Dynodroid (dynodroid) are built on Monkey by adding simple heuristics to reduce wasted inputs. While random testers are scalable and easy to deploy, their lack of app-state awareness often leads to shallow coverage and repeated interactions with the same screens.
Model-based approaches construct an explicit GUI model (e.g., a finite state machine or transition graph) and systematically explore it. Notably, APE (ape) extends Monkey-like random testing with adaptive heuristics to improve coverage and efficiency. However, the models built at runtime are often incomplete or imprecise; they are unable to reason about internal logic and do not support ICC.
Reinforcement learning (RL) has been applied to adaptively guide GUI exploration. ARES (ares) and Q-Testing (qtesting) train models with rewards for novel screens and crashes, thus learning to prioritize unexplored transitions and incrementally improve coverage and bug discovery.
According to recent works (wang2021infrastructure; su2021benchmarking; akinotcho2024mobile), APE is recognized as one of the best-performing GUI testing tools. In subsection 4.2, we compared GAPS against APE, and our empirical evaluation shows that APE was able to reach only 12.82% of the target methods, in contrast to 57.44% achieved by GAPS.
@@ 3.2. LLM UI Automation
A rising trend in Android UI automation is to rely on LLMs to automatically interact with the device screen and complete a task specified by the user. GPTDroid (gptdroid) supplies the LLM with additional contextual information from the view hierarchy, along with a functionality-aware memory that maintains knowledge about past interactions. Autodroid (autodroid) integrates an initial random UI interaction to supply contextual information to the LLM and augment its performance. However, this approach is not exhaustive and scalable with bigger apps. VisionTasker (visiontasker) relies on a different approach by employing an object detection model to find UI elements from screenshots taken from the device. Although this strategy is more resistant to missing or misaligned elements, it is limited by what is currently shown on the screen and cannot properly process dynamic screens (e.g., games) or slow-loading apps. Guardian (guardian) has a replanning strategy that prevents invalid or repeated selection of GUI elements. This approach, coupled with state restoration, allows Guardian to favor app exploration. In subsection 4.2, we evaluated Guardian’s effectiveness in reaching targeted methods within AndroTest applications, achieving an average reachability of only 17.12% over three runs.

### v2

# Mind the GAPS: Bridging the GAPS between Targeted Dynamic Analysis and Static Path Reconstruction in Android Apps

@@ Abstract.
Dynamically executing specific target methods in Android applications remains a critical and unresolved challenge. Despite notable advancements in GUI testing, current tools are insufficient for reliably driving execution toward specific target methods.
To address this challenge, we present GAPS (Graph-based Automated Path Synthesizer), the first system that leverages static, method-guided call graph reconstruction to guide the dynamic, interaction-driven execution of an Android app. GAPS performs a lightweight backward traversal of the call graph, guided by data-flow analysis, to reconstruct paths reaching the target methods. These paths are then translated into instructions that guide runtime app exploration.
On the AndroTest benchmark, GAPS statically identifies paths towards 88.24% of the target methods, averaging just 4.27 seconds per app, and reaching 57.44% of them through dynamic analysis. This performance exceeds the state-of-the-art tools’ one: the model-based GUI tester APE reaches only 12.82%, the hybrid tool GoalExplorer reaches 9.69%, and the LLM-based Guardian reaches 17.12%. Finally, we applied GAPS to the 50 most downloaded apps from the Google Play Store, achieving an average static analysis time of 278.9 seconds to reconstruct paths towards 62.03% of the target methods and reaching 59.86% of them through dynamic analysis.
@@ 1. Introduction
Android dynamic analysis is a methodology that examines an Android app’s behavior during its runtime execution. It can be used to verify security vulnerabilities, measure performance, and detect malicious activities. Complementary to static analysis, which inspects the app without executing it, dynamic analysis involves running the app in a controlled environment, such as a real device, an emulator, while monitoring its interactions with the system, network, and other components. Even though some apps, such as the Android Operating System (OS) services, can come without a Graphical User Interface (GUI), most apps require interaction with the available GUI components to change their status. Thus, dynamic analysis for Android apps mainly involves GUI testing tools. This choice already comes with an important limitation: such tools are able to reach only code within a GUI component and not any code inside an app. In fact, proving the reachability of an app target method is still an open research question (Pradeep et al., 2022; Zeng et al., 2016; Choudhary et al., 2015).
@@
The evolution of Android GUI testing tools has progressed through various phases, each one employing an increasingly sophisticated approach. The very first ones were pseudo-random input generators, such as Monkey (Google, 2025), which lack an understanding of the app’s semantics and UI hierarchy, thus being inefficient for deep functional testing, but very useful for finding shallow random crashes in an app. Then, GUI testers started embedding models, such as finite state machines, to better reproduce common interaction scenarios and maximize exploration. Notably, APE (Gu et al., 2019) is a model-guided GUI testing tool that has been recognized in prior studies for its effectiveness in improving code coverage and detecting GUI-related bugs (Wang et al., 2021; Su et al., 2021; Akinotcho et al., 2025). A hybrid solution, GoalExplorer (Lai and Rubin, 2019), builds a Screen Transition Graph (STG) to statically map relationships between activities and then guides automatic interaction using Stoat (Su et al., 2017a). The advent of machine learning and deep learning techniques paved the way for GUI testers to autonomously discover new test paths, guided by optimization algorithms (Romdhana et al., 2022). Finally, the introduction of Large Language Models (LLMs) enabled GUI testers to leverage semantic and contextual information to solve tasks more closely aligned with those performed by humans. Among recent advancements, Guardian (Ran et al., 2024) stands out for leveraging LLMs to guide UI exploration, showing competitive results across task completion benchmarks. Despite the multiple contributions in the field of Android automatic interaction (Bose et al., 2023; Choudhary et al., 2015; Su et al., 2017b; Mao et al., 2016; Dong et al., 2020; Machiry et al., 2013), the efforts made so far have been mainly driven by either increasing the app code coverage or by replacing a real user in the completion of a task (e.g., “add an event to the calendar app”).
In this paper, we present our solution designed to provide dynamic method-level reachability through the following intuition: every target method should be reached by traversing a path that connects an app entry point to it. Such a path can be statically precomputed and then used to guide the dynamic analysis until the target method is reached.
@@
At first, we evaluated the performance (i.e., the ability to reach a target method) of the best state-of-the-art tools designed for Android GUI interaction and call graph reconstruction. In particular, we randomly selected 50 target methods in each app among the 56 open-source ones from the AndroTest (Choudhary et al., 2015) dataset [^1] https://github.com/samudoria/GAPS/tree/main/evaluation_data/androtest_results, a standard benchmark in the field of Android automatic interaction, still used by current state-of-the-art (Bose et al., 2023; Choudhary et al., 2015; Su et al., 2017b; Mao et al., 2016; Dong et al., 2020; Machiry et al., 2013; Akinotcho et al., 2025). We then tested the performance of the selected tools. We started with the GUI testers, evaluating their ability to dynamically reach the target methods within a 5-minute timeout for each app, over three runs: on average, Guardian (Ran et al., 2024) was able to execute only 17.12% of the target methods, while APE (Gu et al., 2019) and GoalExplorer (Lai and Rubin, 2019) just 12.82% and 9.69% respectively, within the same time limit. In terms of call graph reconstruction, we chose FlowDroid (Arzt et al., 2014), the best static analysis tool for such a purpose (Samhi et al., 2024), and DroidReach (Borzacchiello et al., 2022), the only available static path reconstruction tool: FlowDroid was able to generate paths towards 58.81% of the target methods, DroidReach towards 9.48% of the methods. Our findings confirm the lack of both static and dynamic solutions for reaching a target method in an app.
@@
To this end, we design GAPS (Graph-based Automated Path Synthesizer), a framework relying on static analysis to guide the dynamic execution until a target method is reached in an Android app. Statically, GAPS performs target-oriented, context-sensitive call graph construction via backward traversal, rather than whole-program analysis, restricting its analysis to the subset of the program relevant to a given target method. It uses data-flow analysis techniques, such as points-to-analysis and constant propagation, to statically reconstruct partial execution paths using backward traversal from a target method, resolve conditional statements, and identify triggering UI elements. These results are translated into high-level interaction instructions that dynamically drive the app to execute the desired behavior.
@@
Additionally, we evaluated GAPS in a real-world security oriented scenario: we ran a rule-based static analyzer, called SPECK (SPRITZ, 2025), to identify potential vulnerabilities, and their corresponding method locations, in the 50 most downloaded apps from the Google Play Store. We then assessed the reachability of these methods through GAPS, generating paths towards 62.03% of the target methods in an average of 278.9 seconds  and dynamically reaching 59.86% of them.
@@
  • We design GAPS, a novel approach that leverages static analysis to guide dynamic analysis and ensure method-level reachability in Android apps. We release the framework, evaluation data, and setup in open source [^2] https://github.com/samudoria/GAPS/.
  • We evaluate GAPS on AndroTest, the state-of-the-art dataset for automatic interaction of Android apps. GAPS is able to generate at least one path towards 88.24% of the target methods in 4.27 seconds on average, and to dynamically reach 57.44% of them, without ever exceeding the 5-minute timeout.
  • We used GAPS in a real-world use case scenario to measure the reachability of potentially vulnerable methods in the 50 most downloaded apps from the Google Play Store. With an average static analysis time of 278.9 seconds  we were able to statically compute the paths for 62.03% of the target methods and to dynamically reach 59.86% of them.
@@ 2.1. GUI-Based Test Input Generators
Dynamic test input generators, often called GUI testers, explore Android apps by interacting with their GUIs. Their core functionality relies on repeatedly querying the device for the current view hierarchy, which is retrieved in XML format and lists all the interactable UI elements on the screen, such as buttons, checkboxes, text fields, and menus. Each GUI tester uses this structured representation to decide which element to interact with next. This decision process is guided by a strategy that may be pseudo-random (e.g., Monkey (Google, 2025)), model-based (e.g., APE (Gu et al., 2019)), machine-learning based (e.g., ARES (Romdhana et al., 2022)), or semantic (e.g., Guardian (Ran et al., 2024)).
GUI testers are typically designed to maximize code coverage or detect UI-related bugs. Pseudo-random strategies rapidly issue events for lightweight testing, while model-based and machine-learning based tools interact heuristically with the UI, often guided by internal models of app states or by optimization algorithms, and prioritize breadth and novelty of interactions rather than triggering specific behaviors.
@@
Despite these advances, GUI testers, whether coverage-oriented or goal-driven, are fundamentally limited in their ability to consistently and specifically reach methods within an app, resulting in an inefficient exploration. Moreover, they are unable to reach methods that are not embedded in a GUI component (e.g., a library method).
@@ 2.2. Call Graph Generation
Several tools, such as FlowDroid (Arzt et al., 2014) and Amandroid (Wei et al., 2018), attempt to build inter-procedural call graphs using context- and flow-sensitive analysis. These tools are computationally intensive and face scalability challenges due to their exhaustive exploration of all app edges. For instance, a previous study (Avdiienko et al., 2015) reported that even with 730GB of RAM and 24 hours of analysis time, FlowDroid failed to analyze certain applications. Additionally, Android call graphs are known to suffer from unsoundness, with missing or incorrect edges caused by unresolved dynamic behavior (Samhi et al., 2024). More specifically, many state-changing events for an Android app are not declared in its code, but are instead part of the Android framework, which implicitly invokes callback methods, such as onCreate, onClick (Cao et al., 2015; Bose et al., 2023). In addition, Android apps consist of multiple interacting components (i.e., Activities, Services, and BroadcastReceivers), each governed by its own lifecycle. Accurately modeling these lifecycles and their transitions is critical for generating meaningful call graphs; however, it is also challenging since they involve implicit invocations and asynchronous handlers that cannot be statically resolved. ICC introduces further complexity, as components interact by sending Intent messages that dynamically determine the control flow across components. Without proper modeling of ICC, the resulting call graph is incomplete and may miss valid transitions altogether (Li et al., 2015). Finally, static analysis must also consider conditional logic: even when a path exists between an entry point and a method, it may be guarded by a conditional statement.
Altogether, these challenges make it clear that full call graph construction for Android apps is not only resource-intensive but also fundamentally limited in terms of soundness and precision (Samhi et al., 2024). Moreover, complete call graph construction is not always necessary or even desirable when the analysis goal is to verify the reachability of a specific method. A full call graph introduces the risk of path explosion, particularly in large applications with deeply nested lifecycles, asynchronous flows, and extensive library usage.
@@ 3. Related Work
In this section, we discuss current state-of-the-art research related to GUI testers, call graph generation, and hybrid analysis.
@@ 3.1. GUI Testers
Automated GUI testing for Android has been approached with various techniques, but generally shares the goals of maximizing code coverage and finding UI bugs. Android’s built-in Monkey tool (Google, 2025) is the archetype of random GUI testing. It simply issues pseudo-random UI and system events with minimal configuration, making it extremely easy to use. Variants like SmartMonkey (Haoyin, 2017) and Dynodroid (Machiry et al., 2013) are built on Monkey by adding simple heuristics to reduce wasted inputs. While random testers are scalable and easy to deploy, their lack of app-state awareness often leads to shallow coverage and repeated interactions with the same screens.
Model-based approaches construct an explicit GUI model (e.g., a finite state machine or transition graph) and systematically explore it. Notably, APE (Gu et al., 2019) extends Monkey-like random testing with adaptive heuristics to improve coverage and efficiency. However, the models built at runtime are often incomplete or imprecise; they are unable to reason about internal logic and do not support ICC.
Reinforcement learning (RL) has been applied to adaptively guide GUI exploration. ARES (Romdhana et al., 2022) and Q-Testing (Pan et al., 2020) train models with rewards for novel screens and crashes, thus learning to prioritize unexplored transitions and incrementally improve coverage and bug discovery.
According to recent works (Wang et al., 2021; Su et al., 2021; Akinotcho et al., 2025), APE is recognized as one of the best-performing GUI testing tools. In subsection 4.2, we compared GAPS against APE, and our empirical evaluation shows that APE was able to reach only 12.82% of the target methods, in contrast to 57.44% achieved by GAPS.
@@ 3.2. LLM UI Automation
A rising trend in Android UI automation is to rely on LLMs to automatically interact with the device screen and complete a task specified by the user. GPTDroid (Liu et al., 2024) supplies the LLM with additional contextual information from the view hierarchy, along with a functionality-aware memory that maintains knowledge about past interactions. Autodroid (Wen et al., 2024) incorporates an initial random UI interaction to provide contextual information to the LLM, thereby enhancing its performance. However, this approach is neither exhaustive nor scalable for larger applications. VisionTasker (Song et al., 2024) employs a different approach, utilizing an object detection model to identify UI elements in screenshots taken from the device. Although this strategy is more resistant to missing or misaligned elements, it is limited by what is currently shown on the screen and cannot properly process dynamic screens (e.g., games) or slow-loading apps. Guardian (Ran et al., 2024) has a replanning strategy that prevents invalid or repeated selection of GUI elements. This approach, coupled with state restoration, allows Guardian to favor app exploration. In subsection 4.2, we evaluated Guardian’s effectiveness in reaching targeted methods within AndroTest applications, achieving an average reachability of only 17.12% over three runs.